OpenStudy (anonymous):

Find the value of the given trigonometric function. sec (-300

OpenStudy (anonymous):

sec is inverse cosine so it is cosine 1/-300

OpenStudy (anonymous):

or just use the sec function on your calculator

OpenStudy (anonymous):

i have a t.i30

OpenStudy (anonymous):

\[\sec \left( -\theta \right)=\sec \theta \] \[\sec \left( -300 \right)=\sec 300\] \[\sec 300=\sec \left( 360-60 \right)=\sec 60=\frac{ 1 }{ \cos 60 }=\frac{ 1 }{ \frac{ 1 }{ 2 }}=2 \]